# VocalNet: Speech LLM with Multi-Token Prediction for Faster and High-Quality Generation

### v1

@@ Abstract
Speech large language models (LLMs) have emerged as a prominent research focus in speech processing. We propose VocalNet-1B and VocalNet-8B, a series of high-performance, low-latency speech LLMs enabled by a scalable and model-agnostic training framework for real-time voice interaction. Departing from the conventional next-token prediction (NTP), we introduce multi-token prediction (MTP), a novel approach optimized for speech LLMs that simultaneously improves generation speed and quality. Experiments show that VocalNet outperforms mainstream Omni LLMs despite using significantly less training data, while also surpassing existing open-source speech LLMs by a substantial margin. To support reproducibility and community advancement, we will open-source all model weights, inference code, training data, and framework implementations upon publication.
@@ 3.3 Streaming Speech Decoding
Figure 2: (a) Non-Streaming Attention Mask: $\boldsymbol{v}_{LLM}^{i}$ attends to the complete text positions, and $\boldsymbol{s}^{i}$ attends to the complete text positions and its previous speech positions; (b) Streaming Attention Mask: $\boldsymbol{v}_{LLM}^{i}$ attends to itself and its previous text positions, and $\boldsymbol{s}^{i}$ attends to chunk-limited text positions, itself and its previous speech positions.
@@ 5.2 Model Configuration
We propose VocalNet-1B and VocalNet-8B built upon LLaMA-3.2 1B  [^1] https://huggingface.co/meta-llama/Llama-3.2-1B and LLaMA-3.1 8B  [^2] https://huggingface.co/meta-llama/Llama-3.1-8B respectively. Both models utilize Whisper-large-v3 [27] as the speech encoder and the flow-matching model along with the HiFi-GAN vocoder from CosyVoice 2 to construct the speech response. The downsample adaptor is a 2-layer linear for feature compression with a downsample factor of 5. The speech projector consists of 2-layer Llama decoder layers. The speech decoder comprises 4-layer Llama decoder layers with 2048 hidden size, 32 attention heads, and an 8192-dimensional feed-forward network. Each MTP module is constructed using a single-layer Llama decoder layer with a linear output head. For streaming decoding, the chunk size $C_{s}$ and $C_{t}$ are set to 15 and 5 respectively.

### v2

@@ Abstract
Speech large language models (LLMs) have emerged as a prominent research focus in speech processing. We introduce VocalNet-1B and VocalNet-8B, a series of high-performance, low-latency speech LLMs enabled by a scalable and model-agnostic training framework designed for real-time voice interaction. Central to our contribution is the first application of multi-token prediction (MTP) to speech LLMs. This approach represents a paradigm shift from standard next-token prediction (NTP), offering simultaneous improvements in generation speed and quality. Informed by analysis of MTP’s effect on speech generation and experimental comparisons, we designed a straightforward and highly effective MTP implementation. Experiments demonstrate that VocalNet performs on par with mainstream Omni LLMs even with limited training data, and significantly surpasses existing open-source speech LLMs. To foster reproducibility and community advancement, all model weights, inference code, training data, and framework implementations have been made publicly available at https://github.com/SJTU-OmniAgent/VocalNet.
@@ 3.3 Streaming Speech Decoding
Figure 2: (a) Non-Streaming Attention Mask: $\boldsymbol{v}_{LLM}^{i}$ attends to itself and all text positions, and $\boldsymbol{s}^{i}$ attends to itself, all text positions, and its previous speech positions; (b) Streaming Attention Mask: $\boldsymbol{v}_{LLM}^{i}$ attends to itself and its previous text positions, and $\boldsymbol{s}^{i}$ attends to itself, chunk-limited text positions, and its previous speech positions.
@@ 5.2 Model Configuration
We propose VocalNet-1B and VocalNet-8B built upon LLaMA-3.2-1B-Instruct  [^1] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct and LLaMA-3.1-8B-Instruct  [^2] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct respectively. Both models utilize Whisper-large-v3 [27] as the speech encoder and the flow-matching model along with the HiFi-GAN vocoder from CosyVoice 2 to construct the speech response. The downsample adaptor is a 2-layer linear for feature compression with a downsample factor of 5. The speech projector consists of 2-layer Llama decoder layers. The speech decoder comprises 4-layer Llama decoder layers with 2048 hidden size, 32 attention heads, and an 8192-dimensional feed-forward network. Each MTP module is constructed using a single-layer Llama decoder layer with a linear output head. For streaming decoding, the chunk size $C_{s}$ and $C_{t}$ are set to 15 and 5 respectively.